OpenStudy (anonymous):

Solve the logarithmic equation. log(x+3)=1-log x

OpenStudy (lgbasallote):

isolate all the ones with x first..

OpenStudy (anonymous):

thts wht you'll get: logx+log3=1-logx 2logx=1-log3

OpenStudy (lgbasallote):

logx + log3 = log(3x) according to Product Law. Might wanna recheck that @math456 ^_^

OpenStudy (anonymous):

log 10=1 so log(x+3)=log 10-log x log(x+3)+log x=log10 log(x^2+3x)=log10 get the antilog of both sides to end up with x^2+3=10 x^2+3-10=0 (x+5)(x-2)=0 x=-5 or x=2

OpenStudy (anonymous):

Thank you!